# LLM Agents for Automated Dependency Upgrades

### v1

@@ II Related Work
The notion of automatically updating a codebase to ensure compatibility with newer dependencies has been explored previously, with a focus on identifying differences between library versions. Dagenais and Robillard introduced SemDiff [10], a tool that recommends adaptations to client programs by examining how a framework adjusted to its own modifications. The concept is that if a method is removed from a library, one can pinpoint all callers of that method and assess how they adapted to its removal, thereby proposing a similar solution for the developer’s codebase. Similarly, Nguyen et al. developed LIBSYNC, a tool that assists developers in adapting API usage code by learning complex API usage adaptation patterns from other clients that have already transitioned to a new library version [11].
A key aspect of most of these studies is their focus on statically typed programming languages. However, more recent research has addressed dynamically typed languages. For example, Nielsen et al. aimed to identify locations affected by breaking changes in a codebase using a collection of semantic patches in JavaScript [12]. Various other methodologies for processing code and related documentation have been explored, including manual or rule-based labelling of changelogs, basic NLP techniques like TF-IDF and word embeddings for matching error messages with relevant documentation [13], and popular end-to-end neural approaches.
One dominant approach was neural machine translation (NMT), adapted from NLP, which frames code-related tasks such as bug fixing [14], code summarization [15], or code review automation [16] as sequence-to-sequence prediction tasks using deep learning architectures like encoder-decoder LSTM [17] [18] or Transformer [19]. Despite the versatility of the seq2seq NLP paradigm, its application to source code generation presents unique challenges, such as syntactic validity and global semantic coherence. To address these challenges, extensions like code-aware search strategies [14] and syntax-aware edit-decoders [20] were proposed. Manipulating abstract syntax trees rather than raw code lines has shown promise in better capturing syntax [21]. Alamir et. al showed this by implementing transition-based parsing on ASTs [22] which went on to become a tool [23].
More recent works have used pre-trained Language Models of Code (CodeLMs) such as Liu et. al who used CodeBERT, CodeGPT, CodeT5, amongst others, for code updates. Their findings reveal that CodeLMs, while performing well without temporal constraints, struggle in realistic time-sensitive scenarios, particularly with generalizing to new projects, handling larger methods, and complex updates, often generating null updates [24]. One solution, PCREQ is a 6-component approach evaluated against popular LLMs such as GPT-4o and DeepSeek V3/R1 and outperforms by 18%-20% [25]. Nevertheless, automated program repair tasks have seen success with the use of not only LLMs, but multi-agent LLM systems when testing on the SWE-Bench dataset [26].
As such, in this paper, we address the challenge of code updates by conceptualizing it as a sequence of bug-fixing tasks. Our approach involves a multi-agent LLM system that systematically utilizes code summaries to identify necessary updates, applies them, and iteratively resolves any issues that arise until the code is successfully built and all unit tests pass. This framework also automatically identifies when to hand over to a human for manual intervention.
@@ III Methodology
Our approach (depicted in Fig. 1) makes use of a novel change localisation mechanism called Meta-RAG. We introduced Meta-RAG as a means for efficient and accurate code editing mechanism for large code bases [27]. Meta-RAG works with condensing the codebase by summarising it into a structured natural language replica. This reduces the codebase size in tokens by a large extent (79.9% for SWE-bench Lite dataset [27]) which acts as a preprocessing step to help with information retrieval for change localisation in the codebase by reducing number of tokens to perform this task, as well as providing a more coherent large picture of the codebase. The main process, which is a customised ALMAS framework [9], is following a loop of making changes and resolving issues until the project builds and all tests are passing; the same process is followed by developers to do an upgrade.
@@ III-A Preprocessing
Summary generation is performed as a preprocessing step. Summary agent takes in the repository and builds code summaries and stores them in a data structure that aligns with the code Abstract Syntax Tree (AST). Details of this process could be find in Tawosi et al. [27]. Briefly: LLM agent is asked to summarise each code file into a properly indented file, where each code unit (function or class) gets a singe line summary of its responsibility. Each file also gets a single line summary at file level which describes responsibility or purpure of the code file as a code unit. These LLM generated summaries are matched to the structure of the AST pasted from the code file to ensure correct mapping between the code units in the code file and the summary units in the summary file. These summaries are then traversed to find the relevant pieces of code for edit, thus using RAG on metadata instead of on the code itself. This summarization is done on the entire codebase as a one-off step, and only relevant parts are updated going forward as the codebase is updated.
@@ III-B1 Control Agent
The control agent is another part of the Meta-RAG framework. Given a task, it is responsible for finding the code units “to read” that are useful as a context for a coding agent to perform the task, and the code units that are required to be modified (existing) or authored (new), i.e., “to write”. The context-to-read is to motivate the code agent to reuse the code from the repo and integrate the changes in a natural way to the code base. The control agent uses summaries and also a description of the current task to create instructions for code retrieval. In our solution, the control agent does this by checking file-level summaries first (one line summary per code file in the codebase) to select the relevant files. Then, checking the full summary of the selected files to code units that will form the context. Having the selected code summary units, the counterpart code units can then be fetched from the code files to put together a context for the code agent.
@@ III-B2 Code Agent
The code agent could be any generic autonomous coding agent, in this paper we use GPT-4o. It receives a task and code units as context (to read, and to modify/write) and implements a change. The context is minimised in length by control agent, instead of multiple code search operations commonly used by other code editing agents. This agent implements the code directly into the code files and notifies the summary agent to update the summaries for the modified code, if needed. It is essential to keep the code summaries up-to-date with the codebase.
III-B3 Process Workflow
The upgrade process begins when the user initiates it. The target version can be specified by the user or automatically retrieved from the artifactory to update to the latest available version. The control agent accesses the code base and generates a structure for every pom and yml file in the project, as these files are not included in the summary base (which only summarizes code files). Additionally, the agent consults a migration guide for the target version, which provides general instructions for the upgrade. If a migration guide is unavailable, the control agent will rely on its internal knowledge to perform the upgrade. The migration guide may contain organization-specific changes necessary for the upgrade that are not publicly available, and which the LLM-based agent would not be aware of.
Using the available information, the control agent is tasked with producing initial upgrade instructions, deciding which files need modification and specifying the changes to be applied. The code agent then takes these instructions and implements the changes, which typically involve updating dependencies and altering configurations. After implementing the changes from the initial upgrade instructions, the project is compiled. If the compilation is successful, unit tests are executed. If all tests pass, the upgrade process is deemed successful. However, if the compilation fails or any unit tests fail, the error log is sent to the control agent for resolution. This process resembles Automated Program Repair (APR), where agents are tasked with editing the codebase to fix issues. The cycle of applying changes, compiling the code, running tests, and reviewing the output log continues until the build and test run is successful. The code agent indicates its inability to resolve the issue or if the same error occurs more than $n$ times (with $n=3$), the run stops. These conditions are implemented to prevent an infinite loop that could exhaust resources. This framework also supports collaboration, allowing AI agents to hand over the process to a human developer with a summary of the steps taken. If the developer resolves the issue and determines the process is incomplete, they can re-initiate it and return control to the AI agent to continue.
IV Upgrades Use Case
We evaluate our framework by applying it to upgrade a Java Moneta-based framework for three synthetic repositories. This section describes these repositories and the motivation behind evaluation with synthetic repositories. We also discuss the results of our framework in terms of similarity to the developer upgrade (i.e. Gold), and compare with an open-source agentic software development solution; OpenHands [28].
Figure 2: Illustration of synthetic code repositories used to evaluate the multi-agent LLM framework
@@ IV-A Synthetic Dataset
The dependency to be upgraded is a cloud native event-sourced microservice framework for Java, built upon Spring Boot, and packaged into a dependency. Thousands of code repositories in industry use this framework. However, when an upgrade is made, not all required code updates are applied to all code repositories. As such, in order to test that our framework is able to make all updates covered in a particular upgrade, it would require an infrastructure setup that scales to hundreds of code repositories. In order to overcome this, we have created three synthetic code repositories that cover all changes required for an upgrade (Fig. 2).
The synthetic repositories are upgraded manually to act as a ”Gold” standard by which we compare our solution. As the synthetic repositories represent those we see in a realistic industrial setting, they include many yaml configuration files, pom files, in addition to the Java source code files.
@@ IV-C Discussion
Nonetheless, when looking at precision, LADU overall attempts fewer changes and thus maintains a higher precision, this is important in a realistic setting in order to avoid unwanted code changes in the repository during an upgrade. For instance, in upgrade 3.2 to 3.3, OpenHands removes 128 lines in the codebase, 22 of which are correct (17.2% precision), whilst LADU removes 7 lines, 5 of which were correct (71.4% precision). Thus our framework is able to have a higher confidence in making changes.
Table II results indicate that LADU maintains a significantly reduced number of steps required for the upgrade process, with only 18 steps needed for version 3.1 compared to 106 steps for OpenHands. Consequently, LADU also demonstrates a low execution runtime. Additionally, the framework uses fewer tokens, particularly in version 3.3, where it utilizes 14,387 tokens compared to OpenHands’ 594,366 tokens, due to the Meta-RAG mechanism that condenses the codebase. It is worth noting that the number of tokens used for summary generation are also included in LADU’s token count. Thus our approach is able to efficiently achieve comparable results to state-of-the-art benchmarks.

### v2

@@ II Related Work
Automated codebase updates for dependency compatibility have been explored previously, often by analyzing differences between library versions. SemDiff [10] recommends client code adaptations by tracking framework changes, while LIBSYNC [11] learns API adaptation patterns from clients that have already migrated to newer versions.
Most prior work focuses on statically typed languages, though recent efforts address dynamic languages. For example, Nielsen et al. use semantic patches to locate breaking changes in JavaScript [12]. Other approaches include manual changelog labeling, basic NLP for error-documentation matching [13], and neural methods.
Neural machine translation (NMT) models, adapted from NLP, treat code tasks like bug fixing [14], summarization [15], and review automation [16] as sequence-to-sequence problems using LSTM [17] [18] or Transformer architectures [19]. However, source code generation faces challenges in syntax and semantics, prompting solutions like syntax-aware decoders [20] and AST-based methods [21, 22, 23].
Recent work leverages pre-trained code language models (CodeLMs) such as CodeBERT, CodeGPT, and CodeT5 for code updates. Liu et al. found that CodeLMs struggle with time-sensitive, complex updates and generalization [24]. PCREQ, a six-component system, outperforms popular LLMs by 18–20% [25]. Multi-agent LLM systems have also shown promise in automated program repair on benchmarks like SWE-Bench [26].
Our work frames code updates as sequential bug-fixing tasks, using a multi-agent LLM system [9] that utilises code summaries, iteratively updates code, and ensures successful builds and tests, with automatic handover to human intervention when needed.
@@ III Methodology
Our approach (depicted in Fig. 1) leverages Meta-RAG, a change localisation mechanism designed for efficient code editing in large codebases by summarising code into structured natural language, reducing token count by nearly 80% and enabling scalable information retrieval for localisation tasks [27]. This summarisation step provides a concise, coherent overview of the codebase, facilitating downstream editing and upgrades. The main workflow is built on a customised ALMAS multi-agent framework, which iteratively applies changes and resolves issues until the project builds and all tests pass, mirroring standard developer upgrade practices [9].
@@ III-A Preprocessing
As a preprocessing step, the Summary Agent generates and stores code summaries aligned with the code’s Abstract Syntax Tree (AST). Each code file and its units (functions/classes) receive one-line summaries describing their responsibilities, ensuring accurate mapping between code and summary structures [27]. These summaries are then used for retrieval and localisation, allowing RAG to operate on metadata rather than raw code. Summarisation is performed once for the entire codebase, with incremental updates as the code evolves.
@@ III-B1 Control Agent
The Control Agent, central to Meta-RAG and ALMAS [9], identifies code units to read (for context) and to write (for modification or creation) based on the current task. It first examines file-level summaries to select relevant files, then drills down to finer-grained summaries to pinpoint code units for context and editing. This enables efficient code retrieval and encourages code reuse and natural integration of changes.
@@ III-B2 Code Agent
The Code Agent (e.g., GPT-4o) receives the task and selected code units, then implements the required changes. By minimising context length, the agent avoids repeated code searches typical of other approaches. After editing, it triggers the Summary Agent to update summaries, maintaining alignment between code and metadata
@@
III-B3 Process Workflow
The upgrade process starts with user initiation, specifying a target version or retrieving it automatically from artifactory. The control agent accesses the codebase and generates a structure for every pom and yml file in the project, as these files are not included in the summary base. The Control Agent also consults migration guides if available. It generates upgrade instructions, identifying files and changes needed. The Code Agent applies these changes, updating dependencies and configurations. The project is then compiled and tests ran; if errors occur, logs are sent back to the Control Agent for further resolution, following an Automated Program Repair (APR) loop. The cycle of applying changes, compiling the code, running tests, and reviewing the output log continues until the build and test run is successful. The code agent indicates its inability to resolve the issue or if the same error occurs more than $n$ times (with $n=3$), the run stops. These conditions are implemented to prevent an infinite loop that could exhaust resources. The framework supports collaboration, allowing handover to human developers with a summary of actions taken (stored while the process was running), and resumption by AI agents if needed.
IV Upgrades Use Case
We assess our framework by upgrading a Java Moneta-based system across three synthetic repositories. This section outlines the repositories, explains the rationale for using synthetic data, and presents results comparing our framework’s upgrades to manual developer upgrades (“Gold standard”) and to the OpenHands agentic development tool [28].
@@ IV-A Synthetic Dataset
The target dependency is a cloud-native, event-sourced microservice framework for Java, built on Spring Boot and widely used in industry. Upgrades often require multiple code changes across repositories, but not all updates are consistently applied. To ensure comprehensive evaluation, we created three synthetic repositories that encompass all necessary upgrade changes. These repositories were manually upgraded to serve as the Gold standard for comparison. They reflect real-world industrial setups, including Java source files, yml configurations, and pom files.
@@ IV-C Discussion
In terms of precision, LADU attempts fewer changes, resulting in higher precision and reducing the risk of unwanted modifications. For example, in the 3.2 to 3.3 upgrade, LADU’s removals were more accurate (71.4% vs. OpenHands’ 17.2%).
Table II shows LADU requires far fewer steps (18 vs. 106 for OpenHands) and runs faster. It also uses significantly fewer tokens, especially in version 3.3, thanks to Meta-RAG’s codebase condensation. LADU’s token count includes summary generation, yet it still achieves results comparable to state-of-the-art benchmarks with greater efficiency.